Question 2

Paco chemicals has one product, which requires input from three types of materials to produce batches of soap. Standard cost details for a single batch are shown below:

Material type                      Standard quantity (kgs)                   Standard price per kg ($)

S1                                                8                                                                0.30

S2                                                5                                                                0.50

S3                                                3                                                                0.40

A standard loss of 10% of unit is expected. Actual output was 15408 kgs for the previous week. Details of the material used were:

Material type                        Quantity (kgs)

S1                                             8,284

S2                                              7,535

S3                                              3,334

Required:

  1. Calculate the following material variances;
  1. Cost
  2. Price
  3. Mix
  4. Yield
  1. Explain three significance of analyzing material usage variance into material mix variance and material yield variance
Answer #1
Acutal Output Produced is 15408 kg.The standard quatity of material required for 15408kg of
output is 17120 Kg (15408/90)*100
Statement Showing
"Computation of standard Cost /Actual Cost /Revised actual Quantity
Material Standard Cost Actual Cost Revised Actual Quantity(RAQ KG)
Quantity(SQ)(Kg.) Standard Price(SP) Amount(SQ*SP) Quantity(AQ)(Kg.) Actual Price Amount(AQ*AP)
A 8560 0.3 2568 8284 0.3 2485.2 9577
B 5350 0.5 2675 7535 0.5 3767.5 5985
C 3210 0.4 1284 3334 0.4 1333.6 3591
Total 17120 6527 19153 7586.3 19153
Note:
SQ = Standard Quantity
AQ = Actual Quantity of material consumed
RAQ = Revised Actual Quantity
SP = Standard Price
AP = Actual Price
i) Material Cost Variance
= SQ*SP - AQ*AP =
A = 2568 - 2485.2 = 82.8 F
B = 2675 - 3767.5 = -1092.5 (A)
C = 1284 - 1333.6 = -49.6 (A)
ii) Material Price Variance
= AQ*(SP-AP)
In the current Question Actual price is not given so it has been assumed that Actual price is same as Standard price hence material price variance can not be derived
iii) Material Mix Variance
= SP(RAQ-AQ)
A = 387.75 F
B = -774.844 (A)
C = 102.875 F
iv) Material Yield Variance